# Learning to Learn with Contrastive Meta-Objective

### v3

@@ 4.3 Model Analysis
Here, we show ConML does not require much efforts on tuning hyperparameters. Furthermore, better performance can be obtained through hyperparameter optimization for specific meta-learners. In this Section we show the impact of key ConML settings: (1) the number of subset samples $K$, which influences the model’s complexity, and (2) the contrastive loss, including the distance function $\phi$, the weighting factor $\lambda$, and the use of InfoNCE as a replacement for $(d^{\text{in}}-d^{\text{out}})$.
@@ 4.3.1 Effect of the Number of Subset Samples $K$
Table 4 presents the results from varying the number of subset samples $K$. Starting from $K=1$, we observe moderate performance growth as $K$ increases, while memory usage grows linearly with $K$. Notably, there is a significant discrepancy in both performance and memory (approximately $\sim 2\times$) between the configurations without ConML and with $K=1$. However, $K$ has minimal impact on time efficiency, assuming sufficient memory, since the processes are independent and can be parallelized.
@@ 4.3.2 The Design of Contrastive Loss
Here, we explore various design factors of the contrastive loss. ConML optimizes the following objective: $\mathcal{L}_{\text{ConML}}=\mathcal{L}_{e}+\lambda\mathcal{L}_{c}$. In the previous sections, to highlight our motivation and perform a decoupled analysis, we used the naive contrastive loss $\mathcal{L}_{c}=d^{\text{in}}-d^{\text{out}}$, with the natural cosine distance $\phi(x,y)$. Here, we consider distance function $\phi$ as Euclidean distance, contrastive loss $\mathcal{L}_{c}$ in the form of InfoNCE oord2018representation , varying contrastive weight $\lambda$ in a wide range. More details are provided in Appendix H.1.
@@
4.4 In-Context Learning
@@
Figure 3: Varying the number of in-context examples during inference of In-context learning.
@@
Following (garg2022can, ), we investigate ConML on ICL by learning to learn synthetic functions including linear regression (LR), sparse linear regression (SLR), decision tree (DT) and 2-layer neural network with ReLU activation (NN). We train the GPT-2 (radford2019language, )-like transformer for each function with ICL and ICL w/ ConML respectively and compare the inference (meta-testing) performance. We follow the same model structure, data generation and training settings (garg2022can, ). The detail implementation of ConML is provided in Appendix H.2.
Figure 3 shows the performance, where ICL w/ ConML always makes more accurate predictions than ICL w/o ConML. Table 5 shows the two values to show the effect ConML brings to ICL: Rel. Min. Error is ICL w/ ConML’s minimal inference error given different number of examples, divided by ICL’s; and Shot Spare is when ICL w/ ConML obtain an error no larger than ICL’s minimal error, the difference between the corresponding example numbers. One can observe significant improvement. The effect of ConML on ICL is without loss of generalizability to real-world applications like pre-trained large language models.
@@ 5 Relation with Existing Works
There are existing works attempt to leverage task-identity information into meta-learning, but no "counterpart" for ConML. We discuss them in two categories: (i) The first category is using problem-specific information, while ConML uses problem-agnostic information thus can be plugged-in these methods and brings improvement. They primarily focusing on few-shot image classification problem (doersch2020crosstransformers, ; hiller2022rethinking, ; perera2024discriminative, ). Some require a static pool of base classes for meta-training and class-level alignment (wang2016learning, ; fei2021melr, ; ye2022few, ; tian2022improving, ; wang2023robust, ), while others use image augmentation strategies for sample-level augmentation (li2024metacl, ). These problem-specific approaches are limited by their focus on few-shot classification and cannot effectively handle dynamic or diverse tasks, nor can they generalize to other meta-learning problems beyond classification. As such, they are not directly comparable with ConML. However, ConML can be integrated into these methods. Though they introduce new objectives other than validation loss by additional modules or steps, but they all work under/with the general mini-batch episodic training, either by replacing the steps to obtain $\mathcal{L}_{e}$ in Algorithm 1 with their steps to obtain episodic loss, or introduce additional steps outside Algorithm 1. We demonstrate in Section 4.1 that incorporating ConML leads to performance gains. (ii) The second category includes works that are learner-specific but not problem-specific. For example, gondal2021function and mathieu2021contrastive explore contrastive representations for neural processes. However, their methods are tightly coupled with specific meta-learners that involve explicit model representation vectors, which can be seen as special cases of ConML within amortization-based meta-learners.
@@
Some works include both meta-learning and contrastive learning as key terms but are not directly related to ConML. ni2021close reformulates contrastive learning through meta-learning for better unsupervised learning, while zucchet2022contrastive proposes an optimization-based meta-learner inspired by contrastive Hebbian learning in biology, which is not related to the contrastive learning used in unsupervised learning. lee2023self introduces contrastive set representations for unsupervised meta-learning but does not integrate them with the general meta-learning framework or model. A broader range of related works in meta-learning and contrastive learning are discussed in Appendix I.

### v4

@@ 4.3 Model Analysis
We show ConML does not require much efforts on tuning hyperparameters. Furthermore, better performance can be obtained through hyperparameter optimization for specific meta-learners. In this Section we show the impact of key ConML settings: (1) the number of subset samples $K$, which influences the model’s complexity, and (2) the contrastive loss, including the distance function $\phi$, the weighting factor $\lambda$, and the use of InfoNCE as a replacement for $(d^{\text{in}}-d^{\text{out}})$.
@@ 4.3.1 Effect of the Number of Subset Samples $K$
Table 4 presents the results of varying the number of subset samples $K$. Starting from $K=1$, we observe moderate performance growth as $K$ increases, while memory usage grows linearly with $K$. Notably, there is a significant discrepancy in both performance and memory (approximately $\sim 2\times$) between the configurations without ConML and with $K=1$. However, $K$ has a negligible impact on time efficiency, assuming sufficient memory, as the processes are independent and can be executed in parallel.
@@ 4.3.2 The Design of Contrastive Loss
Here, we explore various design factors of the contrastive loss. ConML optimizes the following objective: $\mathcal{L}_{\text{ConML}}=\mathcal{L}_{e}+\lambda\mathcal{L}_{c}$. In the previous sections, to highlight our motivation and perform a decoupled analysis, we used the naive contrastive loss $\mathcal{L}_{c}=d^{\text{in}}-d^{\text{out}}$, with the natural cosine distance $\phi(x,y)$. Here, we consider distance function $\phi$ as Euclidean distance, contrastive loss $\mathcal{L}_{c}$ in the form of InfoNCE oord2018representation , varying contrastive weight $\lambda$ in a wide range. More details are in Appendix G.1.
@@
In Appendix E, we provide empirical results under synthetic dataset to understand (i) learning to learn with ConML brings generalizable alignment and discrimination abilities; and (ii) alignment enhances fast-adaptation and discrimination enhances task-level generalizability.
4.4 ICL Performance
Following (garg2022can, ), we investigate ConML on ICL by learning to learn synthetic functions including linear regression (LR), sparse linear regression (SLR), decision tree (DT) and 2-layer neural network with ReLU activation (NN). We train the GPT-2 (radford2019language, )-like transformer for each function with ICL and ICL w/ ConML respectively and compare the inference (meta-testing) performance. We follow the same model structure, data generation and training settings (garg2022can, ). More implementation details are provided in Appendix G.2.
Figure 3 shows the performance, where ICL w/ ConML always makes more accurate predictions than ICL w/o ConML. Table 5 shows the two values to show the effect ConML brings to ICL: Rel. Min. Error is ICL w/ ConML’s minimal inference error given different number of examples, divided by ICL’s; and Shot Spare is when ICL w/ ConML obtain an error no larger than ICL’s minimal error, the difference between the corresponding example numbers. One can observe significant improvement. The effect of ConML on ICL is without loss of generalizability to real-world applications like pretrained LLMs.
@@ 4.4 ICL Performance
Figure 3: Varying the number of in-context examples during inference of ICL.
@@ 5 Relation with Existing Works
There are existing works attempt to leverage task-identity information into meta-learning, but no "counterpart" for ConML. We discuss them in two categories: (i) The first category is using problem-specific information, while ConML uses problem-agnostic information thus can be plugged-in these methods and brings improvement. They primarily focusing on few-shot image classification problem (doersch2020crosstransformers, ; hiller2022rethinking, ; perera2024discriminative, ), and require a static pool of base classes for meta-training and class-level alignment (wang2016learning, ; fei2021melr, ; ye2022few, ; tian2022improving, ; wang2023robust, ). These problem-specific approaches are limited by their focus on few-shot classification and cannot effectively handle dynamic or diverse tasks, nor can they generalize to other meta-learning problems beyond classification. As such, they are not directly comparable with ConML. However, ConML can be integrated into these methods. Though they introduce new objectives other than validation loss by additional modules or steps, but they all work under/with the general mini-batch episodic training, either by replacing the steps to obtain $\mathcal{L}_{e}$ in Algorithm 1 with their steps to obtain episodic loss, or introduce additional steps outside Algorithm 1. We demonstrate in Section 4.1 that incorporating ConML leads to performance gains. (ii) The second category includes works that are learner-specific but not problem-specific. For example, gondal2021function and mathieu2021contrastive explore contrastive representations for neural processes. However, their methods are tightly coupled with specific meta-learners that involve explicit model representation vectors, which can be seen as special cases of ConML within amortization-based meta-learners.
@@
Some studies involve both meta-learning and contrastive learning as key components, but they are not directly related to ConML. ni2021close reformulates contrastive learning through meta-learning for better unsupervised learning, while zucchet2022contrastive proposes an optimization-based meta-learner inspired by contrastive Hebbian learning in biology, which is not related to the contrastive learning used in unsupervised learning. lee2023self introduces contrastive set representations for unsupervised meta-learning but does not integrate them with the general meta-learning framework or model.
@@ 6 Conclusion, Limitations and Discussion
Acknowledgment
Y. Wang is sponsored by Beijing Nova Program. Q. Yao is supported by National Natural Science Foundation of China (under Grant No. 92270106) and Beijing Natural Science Foundation (under Grant No. 4242039).
NeurIPS Paper Checklist
  1. 1.
    Claims
  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
  3. Answer: [Yes]
  4. Justification: In this work, we propose ConML, a universal, learner-agnostic contrastive meta-learning framework that emulates the alignment and discrimination capabilities integral to human fast learning, achieved through task-level contrastive learning in the model space. ConML can be seamlessly integrated with meta-training procedure of existing meta-learners, by modifying the conventional mini-batch episodic training, and we provide specific implementations across a wide range of meta-learning algorithms. Empirical results show that ConML consistently and significantly enhances meta-learning performance by improving the meta-learner’s fast-adaptation and task-level generalization abilities. Additionally, we explore in-context learning by reformulating it within the meta-learning paradigm, demonstrating how ConML can be effectively integrated to boost performance.
  5. Guidelines:
    • The answer NA means that the abstract and introduction do not include the claims made in the paper.
    • The abstract and/or introduction should clearly state the claims made, including the contributions made in the paper and important assumptions and limitations. A No or NA answer to this question will not be perceived well by the reviewers.
    • The claims made should match theoretical and experimental results, and reflect how much the results can be expected to generalize to other settings.
    • It is fine to include aspirational goals as motivation as long as it is clear that these goals are not attained by the paper.
  6. 2.
    Limitations
  7. Question: Does the paper discuss the limitations of the work performed by the authors?
  8. Answer: [Yes]
  9. Justification: The paper discusses the limitations of the work performed by the authors, as summarized in the last paragraph of the main text.
  10. Guidelines:
    • The answer NA means that the paper has no limitation while the answer No means that the paper has limitations, but those are not discussed in the paper.
    • The authors are encouraged to create a separate "Limitations" section in their paper.
    • The paper should point out any strong assumptions and how robust the results are to violations of these assumptions (e.g., independence assumptions, noiseless settings, model well-specification, asymptotic approximations only holding locally). The authors should reflect on how these assumptions might be violated in practice and what the implications would be.
    • The authors should reflect on the scope of the claims made, e.g., if the approach was only tested on a few datasets or with a few runs. In general, empirical results often depend on implicit assumptions, which should be articulated.
    • The authors should reflect on the factors that influence the performance of the approach. For example, a facial recognition algorithm may perform poorly when image resolution is low or images are taken in low lighting. Or a speech-to-text system might not be used reliably to provide closed captions for online lectures because it fails to handle technical jargon.
    • The authors should discuss the computational efficiency of the proposed algorithms and how they scale with dataset size.
    • If applicable, the authors should discuss possible limitations of their approach to address problems of privacy and fairness.
    • While the authors might fear that complete honesty about limitations might be used by reviewers as grounds for rejection, a worse outcome might be that reviewers discover limitations that aren’t acknowledged in the paper. The authors should use their best judgment and recognize that individual actions in favor of transparency play an important role in developing norms that preserve the integrity of the community. Reviewers will be specifically instructed to not penalize honesty concerning limitations.
  11. 3.
    Theory assumptions and proofs
  12. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
  13. Answer: [Yes] .
  14. Justification: For each theoretical result, the paper provides the full set of assumptions and a complete (and correct) proof, involving assumption and results in Section 3.2 and proof in Appendix B.
  15. Guidelines:
    • The answer NA means that the paper does not include theoretical results.
    • All the theorems, formulas, and proofs in the paper should be numbered and cross-referenced.
    • All assumptions should be clearly stated or referenced in the statement of any theorems.
    • The proofs can either appear in the main paper or the supplemental material, but if they appear in the supplemental material, the authors are encouraged to provide a short proof sketch to provide intuition.
    • Inversely, any informal proof provided in the core of the paper should be complemented by formal proofs provided in appendix or supplemental material.
    • Theorems and Lemmas that the proof relies upon should be properly referenced.
  16. 4.
    Experimental result reproducibility
  17. Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
  18. Answer: [Yes]
  19. Justification: The experiment setting and methods are fully described.
  20. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • If the paper includes experiments, a No answer to this question will not be perceived well by the reviewers: Making the paper reproducible is important, regardless of whether the code and data are provided or not.
    • If the contribution is a dataset and/or model, the authors should describe the steps taken to make their results reproducible or verifiable.
    • Depending on the contribution, reproducibility can be accomplished in various ways. For example, if the contribution is a novel architecture, describing the architecture fully might suffice, or if the contribution is a specific model and empirical evaluation, it may be necessary to either make it possible for others to replicate the model with the same dataset, or provide access to the model. In general. releasing code and data is often one good way to accomplish this, but reproducibility can also be provided via detailed instructions for how to replicate the results, access to a hosted model (e.g., in the case of a large language model), releasing of a model checkpoint, or other means that are appropriate to the research performed.
    • While NeurIPS does not require releasing code, the conference does require all submissions to provide some reasonable avenue for reproducibility, which may depend on the nature of the contribution. For example
      1. (a)
        If the contribution is primarily a new algorithm, the paper should make it clear how to reproduce that algorithm.
      2. (b)
        If the contribution is primarily a new model architecture, the paper should describe the architecture clearly and fully.
      3. (c)
        If the contribution is a new model (e.g., a large language model), then there should either be a way to access this model for reproducing the results or a way to reproduce the model (e.g., with an open-source dataset or instructions for how to construct the dataset).
      4. (d)
        We recognize that reproducibility may be tricky in some cases, in which case authors are welcome to describe the particular way they provide for reproducibility. In the case of closed-source models, it may be that access to the model is limited in some way (e.g., to registered users), but it should be possible for other researchers to have some path to reproducing or verifying the results.
  21. 5.
    Open access to data and code
  22. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
  23. Answer: [Yes]
  24. Justification: Code is provided at https://github.com/ovo67/ConML_Code.
  25. Guidelines:
    • The answer NA means that paper does not include experiments requiring code.
    • Please see the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • While we encourage the release of code and data, we understand that this might not be possible, so “No” is an acceptable answer. Papers cannot be rejected simply for not including code, unless this is central to the contribution (e.g., for a new open-source benchmark).
    • The instructions should contain the exact command and environment needed to run to reproduce the results. See the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • The authors should provide instructions on data access and preparation, including how to access the raw data, preprocessed data, intermediate data, and generated data, etc.
    • The authors should provide scripts to reproduce all experimental results for the new proposed method and baselines. If only a subset of experiments are reproducible, they should state which ones are omitted from the script and why.
    • At submission time, to preserve anonymity, the authors should release anonymized versions (if applicable).
    • Providing as much information as possible in supplemental material (appended to the paper) is recommended, but including URLs to data and code is permitted.
  26. 6.
    Experimental setting/details
  27. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
  28. Answer: [Yes]
  29. Justification: Most key details are directly mentioned and discussed, while the others can be figured out in the provided code.
  30. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The experimental setting should be presented in the core of the paper to a level of detail that is necessary to appreciate the results and make sense of them.
    • The full details can be provided either with the code, in appendix, or as supplemental material.
  31. 7.
    Experiment statistical significance
  32. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
  33. Answer: [Yes]
  34. Justification: Most results are provided with standard derivation. Exclusions are Table 3 that requires too much source we could not afford to do multiple times, and neglecting for paper space limitation at Table 4.
  35. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The authors should answer "Yes" if the results are accompanied by error bars, confidence intervals, or statistical significance tests, at least for the experiments that support the main claims of the paper.
    • The factors of variability that the error bars are capturing should be clearly stated (for example, train/test split, initialization, random drawing of some parameter, or overall run with given experimental conditions).
    • The method for calculating the error bars should be explained (closed form formula, call to a library function, bootstrap, etc.)
    • The assumptions made should be given (e.g., Normally distributed errors).
    • It should be clear whether the error bar is the standard deviation or the standard error of the mean.
    • It is OK to report 1-sigma error bars, but one should state it. The authors should preferably report a 2-sigma error bar than state that they have a 96% CI, if the hypothesis of Normality of errors is not verified.
    • For asymmetric distributions, the authors should be careful not to show in tables or figures symmetric error bars that would yield results that are out of range (e.g. negative error rates).
    • If error bars are reported in tables or plots, The authors should explain in the text how they were calculated and reference the corresponding figures or tables in the text.
  36. 8.
    Experiments compute resources
  37. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
  38. Answer: [No] ,
  39. Justification: Table 4 has shown the detailed relative consumption comparing the proposed method and standard method.
  40. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The paper should indicate the type of compute workers CPU or GPU, internal cluster, or cloud provider, including relevant memory and storage.
    • The paper should provide the amount of compute required for each of the individual experimental runs as well as estimate the total compute.
    • The paper should disclose whether the full research project required more compute than the experiments reported in the paper (e.g., preliminary or failed experiments that didn’t make it into the paper).
  41. 9.
    Code of ethics
  42. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
  43. Answer: [Yes]
  44. Justification: The research conducted in the paper conforms, in every respect, with the NeurIPS Code of Ethics
  45. Guidelines:
    • The answer NA means that the authors have not reviewed the NeurIPS Code of Ethics.
    • If the authors answer No, they should explain the special circumstances that require a deviation from the Code of Ethics.
    • The authors should make sure to preserve anonymity (e.g., if there is a special consideration due to laws or regulations in their jurisdiction).
  46. 10.
    Broader impacts
  47. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
  48. Answer: [N/A] ,
  49. Justification: There is no societal impact of the work performed.
  50. Guidelines:
    • The answer NA means that there is no societal impact of the work performed.
    • If the authors answer NA or No, they should explain why their work has no societal impact or why the paper does not address societal impact.
    • Examples of negative societal impacts include potential malicious or unintended uses (e.g., disinformation, generating fake profiles, surveillance), fairness considerations (e.g., deployment of technologies that could make decisions that unfairly impact specific groups), privacy considerations, and security considerations.
    • The conference expects that many papers will be foundational research and not tied to particular applications, let alone deployments. However, if there is a direct path to any negative applications, the authors should point it out. For example, it is legitimate to point out that an improvement in the quality of generative models could be used to generate deepfakes for disinformation. On the other hand, it is not needed to point out that a generic algorithm for optimizing neural networks could enable people to train models that generate Deepfakes faster.
    • The authors should consider possible harms that could arise when the technology is being used as intended and functioning correctly, harms that could arise when the technology is being used as intended but gives incorrect results, and harms following from (intentional or unintentional) misuse of the technology.
    • If there are negative societal impacts, the authors could also discuss possible mitigation strategies (e.g., gated release of models, providing defenses in addition to attacks, mechanisms for monitoring misuse, mechanisms to monitor how a system learns from feedback over time, improving the efficiency and accessibility of ML).
  51. 11.
    Safeguards
  52. Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
  53. Answer: [N/A] .
  54. Justification: The paper poses no such risks.
  55. Guidelines:
    • The answer NA means that the paper poses no such risks.
    • Released models that have a high risk for misuse or dual-use should be released with necessary safeguards to allow for controlled use of the model, for example by requiring that users adhere to usage guidelines or restrictions to access the model or implementing safety filters.
    • Datasets that have been scraped from the Internet could pose safety risks. The authors should describe how they avoided releasing unsafe images.
    • We recognize that providing effective safeguards is challenging, and many papers do not require this, but we encourage authors to take this into account and make a best faith effort.
  56. 12.
    Licenses for existing assets
  57. Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
  58. Answer: [Yes]
  59. Justification: The creators or original owners of assets (e.g., code, data, models), used in the paper, are properly credited and are the license and terms of use explicitly mentioned and properly respected.
  60. Guidelines:
    • The answer NA means that the paper does not use existing assets.
    • The authors should cite the original paper that produced the code package or dataset.
    • The authors should state which version of the asset is used and, if possible, include a URL.
    • The name of the license (e.g., CC-BY 4.0) should be included for each asset.
    • For scraped data from a particular source (e.g., website), the copyright and terms of service of that source should be provided.
    • If assets are released, the license, copyright information, and terms of use in the package should be provided. For popular datasets, paperswithcode.com/datasets has curated licenses for some datasets. Their licensing guide can help determine the license of a dataset.
    • For existing datasets that are re-packaged, both the original license and the license of the derived asset (if it has changed) should be provided.
    • If this information is not available online, the authors are encouraged to reach out to the asset’s creators.
  61. 13.
    New assets
  62. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
  63. Answer: [Yes]
  64. Justification: Yes.
  65. Guidelines:
    • The answer NA means that the paper does not release new assets.
    • Researchers should communicate the details of the dataset/code/model as part of their submissions via structured templates. This includes details about training, license, limitations, etc.
    • The paper should discuss whether and how consent was obtained from people whose asset is used.
    • At submission time, remember to anonymize your assets (if applicable). You can either create an anonymized URL or include an anonymized zip file.
  66. 14.
    Crowdsourcing and research with human subjects
  67. Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
  68. Answer: [N/A]
  69. Justification: The paper does not involve crowdsourcing nor research with human subjects.
  70. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Including this information in the supplemental material is fine, but if the main contribution of the paper involves human subjects, then as much detail as possible should be included in the main paper.
    • According to the NeurIPS Code of Ethics, workers involved in data collection, curation, or other labor should be paid at least the minimum wage in the country of the data collector.
  71. 15.
    Institutional review board (IRB) approvals or equivalent for research with human subjects
  72. Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
  73. Answer:[N/A]
  74. Justification: The paper does not involve crowdsourcing nor research with human subjects.
  75. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Depending on the country in which research is conducted, IRB approval (or equivalent) may be required for any human subjects research. If you obtained IRB approval, you should clearly state this in the paper.
    • We recognize that the procedures for this may vary significantly between institutions and locations, and we expect authors to adhere to the NeurIPS Code of Ethics and the guidelines for their institution.
    • For initial submissions, do not include any information that would break anonymity (if applicable), such as the institution conducting the review.
  76. 16.
    Declaration of LLM usage
  77. Question: Does the paper describe the usage of LLMs if it is an important, original, or non-standard component of the core methods in this research? Note that if the LLM is used only for writing, editing, or formatting purposes and does not impact the core methodology, scientific rigorousness, or originality of the research, declaration is not required.
  78. Answer: [N/A]
  79. Justification: The core method development in this research does not involve LLMs as any important, original, or non-standard components.
  80. Guidelines:
    • The answer NA means that the core method development in this research does not involve LLMs as any important, original, or non-standard components.
    • Please refer to our LLM policy (https://neurips.cc/Conferences/2025/LLM) for what should or should not be described.